### v4

# Projected Augmented Waves (PAW) motivated mixed basis sets for small molecules

Abstract
The success behind many pseudopotential methods, such as Projected Augmented Waves (PAW) and Phillips-Kleinman pseudopotential method, is that these methods are all electron methods in disguise. For the Phillips Kleinman and PAW pseudopotential methods we show that there is an explicit all electron reformulation. In the all electron reformulation, in the basis set, there are regular low wavevector basis wave functions and several, specially chosen, high wavevector basis wave functions that are specialized to the atomic environment relevant to nuclei of the substances studied. Using this as motivation, here we propose a new, PAW method motivated, basis set for small molecules, where we use the LO or lo (Localized Orbitals) basis wave functions, and pair them with a Gaussian basis, e.g. Gaussian Type Orbitals (GTO), for a hybrid basis for small molecules. Contracted Gaussian functions are also possible (CGFs). Several different, but related, LO (lo) basis wave functions are considered. We also show how to extend this idea to Slater basis set, Slater type orbitals (STO), instead of GTO orbitals, combined with LO or lo orbitals. This is done using shape functions.
I Introduction
It is of paramount importance to make further progress on the basis set problem for Density Functional Theory (DFT) for small molecules. Small molecules larger then diatomic: carbon dioxide, water, ammonia, methane, ethyne, ethene, ethane to name a few have a variety of uses in industry sciences and engineering [2, 1]. There has been as much progress in optimizing basis sets for small molecule DFT calculations, similarly to the basis sets that have paved breakthroughs in solid state DFT [3, 4, 5, 9, 11, 12, 17, 14, 13, 8, 7, 6, 1, 10, 15, 16] - for crystalline solids. Much progress has been made using ideas based on Linear Combinations of Atomic Orbitals (LCAO) [19, 2, 18], where a good basis for the entire molecule consists of linear combinations of basis functions for the individual atoms. There are many candidate LCAO basis sets:
  • Gaussian basis sets (GTO) based on cartesian Gaussian Type Orbitals (GTO) which are of the form of Gaussians multiplied by polynomials in the cartesian co-ordinates [20, 18, 21, 2, 22, 23]. This is advantageous as four center integrals are efficiently done for this basis set.
  • Slater basis sets where Slater Type Orbitals (STOs) are used in the LCAO method. These are spherical harmonics multiplied by exponentially decaying radial wave functions. It is possible to have multiple wave functions per angular momentum channel - so called multi-zeta basis sets. STO’s have advantages over GTOs as they have the correct exponentially decaying behavior at infinity and the right Kato cusp at the origin [18, 2], however four center integrals are much harder - usually leading to shape function methods.
  • Contracted basis sets (Contracted Gaussian Functions (CGF)) where multiple Gaussians are combined into one basis element used to mimic STO basis elements while retaining the ability to efficiently study four centre integrals [20, 18, 21, 2].
  • Numerical basis sets where one solves the Khon Sham problem numerically and uses these wave functions as basis wave functions.
Furthermore plane waves, outside the LCAO concept, are often used to study periodically arranged arrays of molecules, with achievable distances between the molecules on the order of 10 $nm$, so that interactions between the molecules may be well neglected as a first approximation. The most common electronic structure method for this situation (plane wave basis sets for arrays of molecules) is the Projected Augmented Waves (PAW) [24, 25, 3, 4] method.
Here we focus on the LCAO method. Motivated by pseudopotential methods in DFT we propose a new (LCAO) hybrid basis set. Pseudopotentials, at their core, map a problem into another problem. The exact all electron wavefunction for the system is mapped onto a smooth (pseudized) wavefunction that may well be represented by a small number of plane waves, that is it is not the all electron wavefunction but simplified. Mappings are then made between the pseudized solution and the all electron wavefunction leading to a solution of the Khon Sham (KS) problem. Two of the pre-eminent examples of pseudopotentials are the Phillips Kleinman and Projected Augmented Waves (PAW) pseudopotentials. Here we show that these pseudopotential methods are all electron methods in disguise, where the all electron problem basis set includes smooth plane wave wave functions (with a low cutoff) and specialized wave functions for orbitals near atomic nuclei. Motivated by this observation, and the extensive numerical success of the PAW method [3, 4], we propose new basis sets for small molecules, where localized LO or lo [28, 5, 6, 1] basis wave functions are augmented with Gaussian wave functions - GTOs - (where Gaussians have been extensively used for the problem and various configuration integrals, including for center ones, for them are well known [18, 2, 3] thereby simplifying the calculation of the four center problem for Coulomb interactions) or STO’s for enhanced accuracy (but with shape functions). This construction is done in a manner similar to Phillips-Kleinman and PAW all electron basis sets when represented as all electron basis sets (see Section II). Multiple types of LO (lo) basis wave functions are possible. Thereby we propose that this new hybrid basis set is efficient for DFT and Hartree-Fock (HF) calculations of small molecules.
II Many pseudopotential methods are all electron methods in disguise (motivation)
In this section we reformulate PAW and the Phillips-Kleinman method as all electron methods. We show the basis set of these all electron methods is composed of plane waves and special wave functions that are chosen to represent states near the core (near the nuclei). We show the two methods to be equivalent in their all electron formulation - depending on the exact choice of the localized wave functions near the nuclei (cores).
II.1 Phillips-Kleinman as an all electron method in disguise
We now recall the Orthogonalized Plane Wave (OPW) wave functions used for the Phillips-Kleinman method. We know that [6, 3]:
$\left|\phi_{OPW}\left(\mathbf{k}+\mathbf{K}\right)\right\rangle$
$=\left|\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\rangle$
$-\sum_{i}\left\langle\phi_{i}\left(\mathbf{r}\right)\mid\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\rangle\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle$ (1)
Here $\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle$ are localized wave functions relevant to the single site problem and $\mu$ labels the atoms. Here $V$ is the volume of the primitive lattice cell and $\mathbf{k}$ is a wavevector in the first Brillouin zone while $\mathbf{K}$ is a reciprocal lattice vector. Now the main claim of the Phillips-Kleinman method (greatly numerically supported [6, 3]) is that for a reasonable number of plane waves the exact Khon Sham (KS) wavefunction $\left|\psi_{n}\right\rangle$ can be written as
$\left|\psi_{n}\right\rangle\cong\sum_{\mathbf{K}}c_{\mathbf{k}}^{\mathbf{K}}\left|\phi_{OPW}\left(\mathbf{k}+\mathbf{K}\right)\right\rangle$ (2)
for some reasonable cutoffs $\mathbf{K}_{max}$ that means that
$\left|\psi_{n}\right\rangle$ $\in Span\left\{\left|\phi_{OPW}\left(\mathbf{k}+\mathbf{K}\right)\right\rangle\right\}$
$\subset Span\left\{\left\{\left|\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\rangle\right\},\left\{\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle\right\}\right\}$ (3)
As such in some situations its better and in many ways simpler to do an all electron calculation with the basis being given by:
$\left\{\left\{\left|\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\rangle\right\},\left\{\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle\right\}\right\}$ (4)
As such Phillips-Kleinman is an all electron method in disguise, this will motivate our basis set constructions in Section III. We shall see below that PAW is very similar.
II.2 PAW is an all electron method in disguise
II.2.1 PAW review
In the PAW formulation of the KS problem, we study the highly oscillatory eigenwavefunction of the KS problem near the nucleus of an atom using plane waves. In order to obtain a low cutoff (which can be as low as ~60 Rydberg in energy) and as such a tolerable basis set we must have that the wavefunction we represent using plane waves (the pseudized wavefunction) is very smooth. The key idea of PAW is then to introduce:
$\left|\psi_{n}\right\rangle=\hat{\mathcal{T}}\left|\tilde{\psi}_{n}\right\rangle$ (5)
where $\hat{\mathcal{T}}$ is some linear transformation (specified below) and $\left|\psi_{n}\right\rangle$ is the exact Khon Sham wave functions (solutions of the KS equations) while $\left|\tilde{\psi}_{n}\right\rangle$ are the pseudized wave functions we can approximate efficiently using plane waves at low cutoffs. Now we construct a set of augmentation spheres and we wish for $\hat{\mathcal{T}}$ to be the identity outside the augmentation spheres (centered around nuclei) , so we write:
$\hat{\mathcal{T}}=\mathbb{I}+\sum_{\mu}S_{\mathbf{\mu}}$ (6)
Here $\mu$ are the atomic augmentation spheres. Now we want the pseudized wave functions $\left|\tilde{\psi}_{n}\right\rangle$ to be smooth inside the sphere and we know that the exact KS wavefunction can be reasonably efficiently approximated by a single atom, atomic wavefunction $\left|\phi_{i\mu}\right\rangle$ near the atomic nuclei. As such if we choose $S_{\mathbf{\mu}}$ to be
$S_{\mu}\left|\tilde{\psi}_{n}\right\rangle=\sum c_{i\mu}^{n}\left|\phi_{i\mu}\right\rangle-\sum c_{i\mu}^{n}\left|\tilde{\phi}_{i\mu}\right\rangle$ (7)
where $\left|\tilde{\phi}_{i\mu}\right\rangle$ are smooth wave functions. Where we choose
$c_{i\mu}^{n}=\left\langle\tilde{p}_{i\mu}\mid\tilde{\psi}_{n}\right\rangle$ (8)
Here $\left|\tilde{p}_{i\mu}\right\rangle$ are the projector wave functions. As such we have that:
$\left|\psi_{n}\right\rangle=\hat{\mathcal{T}}\left|\tilde{\psi}_{n}\right\rangle=\left|\tilde{\psi}_{n}\right\rangle+\sum_{i}\left(\left|\phi_{i\mu}\right\rangle-\left|\tilde{\phi}_{i\mu}\right\rangle\right)\left\langle\tilde{p}_{i\mu}\mid\tilde{\psi}_{n}\right\rangle$ (9)
Furthermore the pseudized wave functions $\left|\tilde{\psi}_{n}\right\rangle$ may be well represented by plane waves.
II.2.2 Reformulation of PAW as an all electron method
The main claim of PAW (which has been extensively numerically verified) is that there is a good expansion in plane waves for the PAW wavefunction given by:
$\left|\psi_{n}\right\rangle\cong\left|\tilde{\psi}_{n}\right\rangle+\sum_{i}\left(\left|\phi_{i\mu}\right\rangle-\left|\tilde{\phi}_{i\mu}\right\rangle\right)\left\langle\tilde{p}_{i\mu}\mid\tilde{\psi}_{n}\right\rangle.$ (10)
That is the wavefunction $\left|\tilde{\psi}_{n}\right\rangle$ is made of a reasonably small number of plane waves of the form $\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)$ in particular we have that:
$\left|\psi_{n}\right\rangle\in Span\left\{\left\{\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\},\left\{\left|\phi_{i\mu}\right\rangle\right\},\left\{\left|\tilde{\phi}_{i\mu}\right\rangle\right\}\right\}$ (11)
to significant accuracy, for small number of plane waves here $\left|\psi_{n}\right\rangle$ are all electron wave functions. Now because $\left|\tilde{\phi}_{i}\right\rangle$ is very smooth we have that
$\left|\tilde{\phi}_{i\mu}\right\rangle\in Span\left\{\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\}$ (12)
As such the main claim of PAW can be written as
$\left|\psi_{n}\right\rangle\in Span\left\{\left\{\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\},\left\{\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle\right\}\right\}$ (13)
for a reasonable number of plane waves. Since the basis set of PAW and Phillips-Kleinman when formulated as all electron methods are similar, both methods should yield, in their all electron formulation, similar accuracy and require basis sets of the same size. They are equivalent within the all electron formulation when they use the same localized orbitals $\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle$.
III GLO, SLO and PLO basis sets
Here we would like to work backwards, we wish to formulate an all electron method as the inverse of a pseudopotential method where we use specialized wave functions (with high wave numbers) near the nuclei and many regular, smooth, wave functions for the whole space. We focus on small molecules and crystalline solids. Below we will simply describe the final answer - that is relevant basis sets.
III.1 Glo and GLO for small molecules
Here G in Glo and GLO stands for Gaussians while lo and LO for localized orbitals and the two types of elements (Gaussians and LO (lo) basis wave functions) are combined into one basis set. We consider the lo basis set [28]:
$\Phi_{lo}^{\mu Elm}=u_{l}^{E_{l}}\left(\left|\mathbf{r}-\mathbf{r}_{\mu}\right|\right)\Theta\left(S_{l}^{\mu}-\left|\mathbf{r-r}_{\mu}\right|\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (14)
where $S_{l}^{\mu}$ is chosen to make the wavefunction continuous everywhere [28]. Here:
$\left[-\frac{d^{2}}{dr^{2}}+\frac{l\left(l+1\right)}{r^{2}}+\bar{V}_{KS}\left(r\right)\right]ru_{l\mu}^{E_{l}}\left(r\right)=E_{l}ru_{l\mu}^{E_{l}}\left(r\right)$ (15)
and $\bar{V}_{KS}\left(r\right)$ is the spherically average KS potential or HF single particle potential. Furthermore $E_{l}$ is chosen at the binding energy of the single atom system (other options are possible for example one could use the eigen-energies from the previous iteration to self consistency in the KS problem [26, 27] for the energies $E_{l}$). Alternatively we can use LO basis sets [28]:
$\Phi_{LO}^{\mu Elm}$ $=u_{l}^{E_{l}}\left(\left|\mathbf{r}-\mathbf{r}_{\mu}\right|\right)\Theta\left(S_{l}^{\mu}-\left|\mathbf{r-r}_{\mu}\right|\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$
$+B_{l}^{\mu}\dot{u}_{l}^{E_{l}}\left(\left|\mathbf{r}-\mathbf{r}_{\mu}\right|\right)\Theta\left(S_{l}^{\mu}-\left|\mathbf{r-r}_{\mu}\right|\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (16)
where $B_{l}^{\mu}$ and $S_{l}^{\mu}$ are chosen to make the wavefunction continuous and continuously differentiable everywhere [28]. Here $\dot{u}_{l}^{E_{l}}=\frac{\partial}{\partial E}u_{l}^{E_{l}}$. Higher derivative terms are also possible [28], also regular LO and lo is possible [5, 6, 1] as well as atom DFT based wave functions that is solutions to the radially averaged KS problem for a single atom [3]. Based on our experience with PAW a very small number of wave functions such as in Eq. (16) will suffice. We also add a large number of Gaussians to our basis set:
$G_{l,m,\sigma}^{\mu}=\frac{1}{\mathcal{N}_{l}}\exp\left(-\frac{\left|\mathbf{r-r}_{\mu}\right|^{2}}{2\sigma}\right)\left|\mathbf{r-r}_{\mu}\right|^{l}Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (17)
to the molecular basis set. The wave functions $G_{l,m,\sigma}^{\mu}$, are smooth and replace the plane waves we have seen before in all electron formulations of pseudopotential methods. Here $\mathcal{N}_{l}$ is a normalization constant and cartesian Gaussians are also possible [20, 18]. This is a basis set for many small molecules. Furthermore Gaussians Coulomb integrals may be efficiently computed, see [30, 18] and Appendix A so there are “few” Coulomb integrals involving complicated functions [31]. Furthermore Eqs. (32) may be used, in the case when there is only one difficult to handle LO (lo) wavefunction, on two of the remaining Gaussians thereby simplifying that case to a three center problem. Overall if we consider the number of LO or lo orbitals as order one (highly reasonable for small molecules based on or experience with PAW [24, 25, 3]) this becomes a three center problem at worst. We note that contracted Gaussian Functions (CGF) are also possible [29, 18].
III.2 Slo and SLO basis sets
We now consider a second hybrid basis set made of Slater type orbitals and localized orbitals. We first consider basis functions of the form given in Eqs. (14) or (16). We augment them with basis wave functions of the form:
$S_{l,m,\xi}^{\mu}=\frac{1}{\mathcal{N}_{l}}\exp\left(-\xi\left|\mathbf{r-r}_{\mu}\right|\right)\left|\mathbf{r-r}_{\mu}\right|^{l}Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (18)
While four center integrals are very difficult with this basis set three center integrals are numerically manageable [29]. Below we describe practical methods to achieve accurate three center decompositions into shape functions.
III.2.1 Practical methods for implementation (three center decompositions)
We consider standard shape functions $\omega_{\kappa}$ for the density and add to them the following additional shape functions:
$\omega_{\mu Elm}^{E^{\prime}l^{\prime}m^{\prime}}\left(\mathbf{r}\right)=\Phi_{lo/LO}^{\mu E^{\prime}l^{\prime}m^{\prime}*}\left(\mathbf{r}\right)\cdot\Phi_{lo/LO}^{\mu Elm}\left(\mathbf{r}\right)$ (19)
this allows for hybridization between the atomic orbitals such as for example the $sp^{3}$ hybridization common to Carbon (if this effect is not important for the molecule just the diagonal terms suffice). Furthermore for notational reasons we absorb the label $\mu;Elm;E^{\prime}l^{\prime}m^{\prime}$ into $\kappa$. We introduce the projectors $P_{\lambda}$ such that
$\int d^{3}\mathbf{r}P_{\lambda}^{*}\left(\mathbf{r}\right)\omega_{\kappa}\left(\mathbf{r}\right)=\delta_{\lambda\kappa}$ (20)
Where for example we may choose:
$M_{\alpha\beta}$ $=\int d^{3}\mathbf{r}\omega_{\alpha}^{*}\left(\mathbf{r}\right)\omega_{\beta}\left(\mathbf{r}\right)$
$P_{\lambda}\left(\mathbf{r}\right)$ $=\sum_{\kappa}\left[M^{-1}\right]_{\lambda\kappa}\omega_{\kappa}\left(\mathbf{r}\right)$ (21)
Then if we introduce the single particle density $\rho\left(\mathbf{r}\right)$ we may write that [29]
$\rho\left(\mathbf{r}\right)$ $\cong\sum_{\kappa}\left[\int d^{3}\mathbf{r}^{\prime}\rho\left(\mathbf{r}^{\prime}\right)P_{\kappa}^{*}\left(\mathbf{r}^{\prime}\right)\right]\omega_{\kappa}\left(\mathbf{r}\right)\equiv\sum_{\kappa}c_{\kappa}\omega_{\kappa}\left(\mathbf{r}\right)$
$c_{\kappa}$ $=\int d^{3}\mathbf{r}^{\prime}\rho\left(\mathbf{r}^{\prime}\right)P_{\kappa}\left(\mathbf{r}^{\prime}\right)$ (22)
We now introduce more notation for the single particle basis set where $\left\{\nu\right\}=\left\{\mu,Elm\right\}\cup\left\{\mu,l,m,\xi\right\}$ and write the eigenstates are given by
$\left|\lambda_{n}\right\rangle=\sum_{\nu}\lambda_{n\nu}\left|\psi_{\nu}\right\rangle$ (23)
We may now introduce [29]:
$P_{\mu\nu}=\sum_{n}f\left(\varepsilon_{n}\right)\lambda_{n\mu}^{*}\lambda_{n\nu}$ (24)
Then the density is given by:
$\rho\left(\mathbf{r}\right)$ $=\sum_{\mu\nu}P_{\mu\nu}\psi_{\mu}^{*}\left(\mathbf{r}\right)\psi_{\nu}\left(\mathbf{r}\right)$
$\rho\left(\mathbf{r}\right)$ $\cong\sum_{\kappa}\sum_{\mu\nu}P_{\mu\nu}\left[\int d^{3}\mathbf{r}^{\prime}\psi_{\mu}^{*}\left(\mathbf{r}^{\prime}\right)\psi_{\nu}\left(\mathbf{r}^{\prime}\right)P_{\kappa}^{*}\left(\mathbf{r}^{\prime}\right)\right]\omega_{\kappa}\left(\mathbf{r}\right)$ (25)
Which is a three center integral. Furthermore the KS equation is now given by:
$H_{\alpha\beta}^{KS}$ $=\left\langle\psi_{\alpha}\mid\frac{-\nabla^{2}}{2m}-\sum_{\mu}\frac{Z_{\mu}e^{2}}{\left|\mathbf{R}_{\mu}-\mathbf{r}\right|}+V_{XC}\left(\mathbf{r}\right)\mid\psi_{\beta}\right\rangle$
$+\sum_{\kappa}c_{\kappa}\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\omega_{\kappa}\left(\mathbf{r}_{1}\right)\frac{e^{2}}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}\psi_{\alpha}^{*}\left(\mathbf{r}_{2}\right)\psi_{\beta}\left(\mathbf{r}_{2}\right)$ (26)
Which is another three center integral. As such it is possible to do practical calculations for small molecules within DFT for SLO and Slo basis sets.
III.3 Plo and PLO basis sets for crystalline solids
Here P in Plo and PLO stands for plane wave while lo and LO stands for localized orbitals. We first consider basis functions of the form given in Eqs. (14) or (16) . We also add plane wave
$\chi_{\mathbf{k}+\mathbf{K}}=\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)$ (27)
to the solid state basis set. This is a basis set for many crystalline solids. However there are many competitor basis sets [9, 11, 12, 13, 14, 17, 3, 8, 7, 5, 6, 1, 10, 15, 16].
IV Conclusions
In this work - motivated by the fact that many pseudopotential methods are all electron methods in disguise with basis sets of the form of slowly oscillating basis wave functions and special high momentum basis wave functions to account for the environment of the nucleus - we proposed several related new basis sets for small molecules and crystalline solids. In this basis set we use the LO or lo wave functions introduced recently in [28] (although regular LO or lo basis wave functions will also do quite well [5, 6, 1]) combined either with Gaussians (GTOs and CGFs) or Slater basis sets - STO - (for molecules) or plane waves (for solids) to obtain a total basis for the system. This would allow for initial calculations for many molecules as the four center calculation is done for Gaussians and two center integrals have been done for very generic functions [2, 31, 18] and for Slater type orbitals it is possible to introduce appropriate density shape functions (which we extended). This should help open small molecules for theoretical DFT or HF like calculations and explorations.
Appendix A Gaussian basis sets (review)
We are interested in four center integrals for the Coulomb problem [20, 18]:
$I_{\alpha\beta\gamma\delta}=$
$\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\frac{1}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}\varphi_{\alpha}^{*}\left(\mathbf{r}_{1}\right)\varphi_{\beta}\left(\mathbf{r}_{1}\right)\varphi_{\gamma}^{*}\left(\mathbf{r}_{2}\right)\varphi_{\delta}\left(\mathbf{r}_{2}\right)$ (28)
This problem scales as $N^{4}$ (where $N$ is the size of the basis set) and must be evaluated for non-crystalline materials (molecules) where the Weinert method does not directly apply. As such it is often the bottleneck for many calculations [20, 18]. We can introduce spherical Gaussians to simplify these integrals:
$G_{lm,n,\mathbf{R},p}\left(\mathbf{r}\right)=Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{R}}\right)\exp\left(-p\left(\mathbf{r}-\mathbf{R}\right)^{2}\right)\left|\mathbf{r}-\mathbf{R}\right|^{2n+l}$ (29)
We can also introduce cartesian Gaussians:
$G_{a,b,c,\mathbf{R},p}\left(\mathbf{r}\right)$
$=\exp\left(-p\left(\mathbf{r}-\mathbf{R}\right)^{2}\right)\left(x-\mathbf{R}_{x}\right)^{a}\left(y-\mathbf{R}_{y}\right)^{b}\left(z-\mathbf{R}_{z}\right)^{c}$ (30)
We now have that:
$G_{lm,n,\mathbf{R},p}\left(\mathbf{r}\right)=\sum_{a+b+c=l+2n}A_{abc}^{lmn}G_{a,b,c,\mathbf{R},p}\left(\mathbf{r}\right),$ (31)
so we might as well focus on cartesian Gaussians for the wave functions in the integral in Eq. (28). Now we have that:
$G_{a,b,c,\mathbf{R},p}\left(\mathbf{r}\right)G_{a^{\prime},b^{\prime},c^{\prime},\mathbf{R}^{\prime},p^{\prime}}\left(\mathbf{r}\right)$
$=\sum_{ABC}P_{ABC}^{abc,\mathbf{R},\mathbf{R}^{\prime}}G_{ABC\bar{\mathbf{R}},p+p^{\prime}}\left(\mathbf{r}\right)$ (32)
Where
$\bar{\mathbf{R}}=\frac{p\mathbf{R}+p^{\prime}\mathbf{R}^{\prime}}{p+p^{\prime}}$ (33)
As such for cartesian Gaussians, up to linear sums, we may as well focus on two center integrals of the form:
$I_{\alpha\beta\gamma\delta}=$
$=\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\frac{1}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}G_{abc,\mathbf{R},p}\left(\mathbf{r}_{1}\right)G_{a^{\prime}b^{\prime}c^{\prime},\mathbf{R}^{\prime},p^{\prime}}\left(\mathbf{r}_{2}\right)$
$=\sum\kappa_{abc}^{ABC}\frac{\partial^{A+B+C}}{\partial\mathbf{R}_{x}^{A}\partial\mathbf{R}_{y}^{B}\partial\mathbf{R}_{z}^{C}}\kappa_{a^{\prime}b^{\prime}c^{\prime}}^{A^{\prime}B^{\prime}C^{\prime}}\frac{\partial^{A^{\prime}+B^{\prime}+C^{\prime}}}{\partial\mathbf{R}^{\prime}{}_{x}^{A^{\prime}}\partial\mathbf{R}^{\prime}{}_{y}^{B^{\prime}}\partial\mathbf{R}^{\prime}{}_{z}^{C^{\prime}}}\times$
×$\times\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\frac{1}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}G_{000,\mathbf{R},p}\left(\mathbf{r}_{1}\right)G_{000,\mathbf{R}^{\prime},p^{\prime}}\left(\mathbf{r}_{2}\right)$ (34)
We now introduce the Boys’ function:
[…]
Then we have that:
$\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\frac{1}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}G_{000,\mathbf{R},p}\left(\mathbf{r}_{1}\right)G_{000,\mathbf{R}^{\prime},p^{\prime}}\left(\mathbf{r}_{2}\right)$
$=\sqrt{\frac{4\alpha}{\pi}}F_{0}\left(\alpha\left|\mathbf{R}-\mathbf{R}^{\prime}\right|^{2}\right)$ (36)
With:
$\alpha=\frac{p\cdot p^{\prime}}{p+p^{\prime}}$ (37)
As such four center integrals may be efficiently evaluated with Gaussian wave functions.
?refname?
  • 1 [2000] E. Sjostedt, L. Nordstrom, and D. J. Singh, Sol. Sta. Comm. 114, 15 (2000).
  • 2 [2019] S. M. Blinder and D. E. Huse, Mathematical Physics in Theoretical Chemistry, (Elsevier Inc. Amsterdam, 2019).
  • 3 [2020] R. M. Martin, Electronic Structures Basic Theory and Practical Methods (Cambridge University Press, Cambridge, 2020).
  • 4 [2009] D. Marx and J. Hutter, Ab Initio Molecular Dynamics Basic Theory and Advanced Methods (Cambridge University Press, Cambridge, 2009).
  • 5 [1991] D. J. Singh, Phys. Rev. B 43, 6388 (1991).
  • 6 [2006] D. J. Singh and D. Nordstrom, Planewaves, pseudopotentials, and the LAPW method (Springer, New York, 2006).
  • 7 [2014] G. Michalicek, Extending the precision and efficiency of all-electron full-potential linearized augment plane-wave density functional theory (Aachen University, 2014, thesis).
  • 8 [2013] G. Michalicek, M. Betzinger, C. Freidrich and S. Blugel, Comp. Phys. Comm. 184, 2670 (2013).
  • 9 [1975] O. K. Andersen, Phys. Rev. B 12, 3060 (1975).
  • 10 [1984] H. L. Skriver, The LMTO method Muffin-Tin Orbitals and Electronic Structure (Springer, New York, 1984).
  • 11 [1984] O. K. Andersen, and O. Jepsen, Phys. Rev. Lett. 53, 2571 (1984).
  • 12 [2003] O. K. Andersen, T. S.-Dasgupta, and S. Ezhof, Bull. Mat. Sci. 26, 19 (2003).
  • 13 [1954] W. Kohn, N. Rostoker, Phys. Rev. 94, $1$ (1954).
  • 14 [1947] J. Korringa, Physica 13, 392 (1947).
  • 15 [1989] J. M. Soler, and A. R. Williams, Phys. Rev. B 40, 1560 (1989).
  • 16 [1990] J. M. Soler and A. R. Williams, Phys. Rev. B 42, 9728 (1990).
  • 17 [1967] T. L. Loucks, Augmented Plane Wave Method (W. A. Benjamin Inc., New York, 1967).
  • 18 [1975] T. Helgaker, P. Jorgensen, J. Olsen, Molecular Electronic Structure Theory (John Wiley & Sons LTD. New York, 2000).
  • 19 [2014] I. N. Levine, Quantum Chemistry (Pearson Education Inc., Boston, 2014).
  • 20 [1995] T. Helgakar and P. R. Taylor, “Gaussian Basis sets and molecular integrals” in Modern Electronic structure theory part II, D. R. Yarkony (eds.) (World Scientific, Singapore, 1995).
  • 21 [1990] D. Feller and E. R. Davidson, Rev. Comp. Chem. 1, 1 (1990).
  • 22 [1972] W. J. Hehre, R. Ditchfield, J. A. Pople, J. Chem. Phys. 56, 2257 (1972).
  • 23 [1973] P. C. Hariharan and J. A. Pople, Theor. Chem. Acta. 28, 213 (1973).
  • 24 [1994] P. E. Blochl, Phys. Rev. B 50, 17953 (1994).
  • 25 [1996] G. Kresse, J. Furthmuller, Phys. Rev. B 54, 11169 (1996).
  • 26 [2024] G. Goldstein, arXiv 2405.11926
  • 27 [2024] G. Goldstein, arXiv 2407.15299
  • 28 [2024] G. Goldstein, arXiv 2403.15954.
  • 29 [2001] W. Koch, and M. C. Holthausen, A chemist’s guide to density functional theory (Wiley, new York, 2001).
  • 30 [1950] S. F. Boys, Proc. Roy. Soc. A200, 542 (1950).
  • 31 [2024] G. Goldstein, arXiv 2404.08493